Circular No. 8258

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

QUASAR OUTBURST IN MENSA A. Clocchiatti and D. Minitti, Pontificia Universidad Católica de Chile,

on behalf of the SuperMACHO project (involving also C. Aguilera, A. Becker, K. Cook, R. Covarrubias, S. Hawley, R. Hiriart, S. Keller, A. Miceli, G. Miknaitis, S. Nikolaev, K. Olsen, J. Prieto, G. Proctor, A. Rest, B. Schmidt, C. Smith, C. Stubbs, N. Suntzeff, and D. Welch), report the discovery of a highly unusual outburst of a quasistellar object (QSO) at z = 2.87. The object, located at α = 5^{h}13^{m}03^{s}.94, δ = −70^{o}22′50′′.5 (equinox 2000.0), had a quiescent broadband magnitude of VR ≈ 22.2. Multiple broadband-VR-filter images taken at the Blanco 4-m telescope at Cerro Tololo reveal no significant change in brightness from the beginning of observations in 2001 Sept. until 2003 Sept. 27 UT, around the date that it began to brighten linearly in flux with time. By Dec. 15, it had brightened by a factor of 15 and has reached VR = 19.4 with no evidence of a change in trend. Identified as a SuperMACHO microlensing candidate, follow-up spectroscopy (456-s exposure) was obtained at the Magellan II (Clay) 6.5-m telescope (+ LDSS2 spectrograph with medium-blue grism) on Dec. 3.8. The spectrum reveals several prominent QSO lines with strong P-Cyg profiles, indicative of an ejection velocity of 1600 km/s; identified features include H (Lyα), N V, C IV 154.9-nm, Si IV 139.3-nm, O IV, and [C III] 191.0-nm. The high redshift, extreme current luminosity, evidence of outflow (from P-Cyg line profiles), and continued brightening suggest that this is an object worthy of intense monitoring.

SUPERNOVAE 2003kw AND 2003kx T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2003kw (cf. IAUC 8252), obtained by M. Calkins on Dec. 17.46 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova. Adopting the NASA/IPAC Extragalactic Database (NED) recession velocity of 7880 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 6300 km/s. A spectrum of SN 2003kx (cf. IAUC 8254), obtained on Dec. 17.50, shows it to be a type-II supernova. Adopting the NED recession velocity of 1860 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 8200 km/s.

2003 December 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green